After using a log-log sales model to analyze the sales data, the marketing department obtains the following own- and cross-price elasticities result.

Change in the price of

Bob’s

Chan’s

Ace

Bob’s

-1.2

.8

.3

Chan’s

.4

-4.2

.6

Ace

.15

1.0

-3.2

Change in

Sales         

Question: The average price of Bob’s new line of cellular devices is about $25,000 with a 40% gross margin after COG. Is it profitable for Bob’s to drop the average price to $20,000?
Answer #1

Change in price= 5000/25000= 20%

So change in demand = 1.2*20%= 24%

Let initial demand be X

So initial profitability= X*25000*0.6= 15000X

New profitability= 1.24*20000*0.6= 14880X

So it is not wise or profitable to reduce price by 20%